# references.bib

### v1

@@ 2.5 A Practical Example: Apple Intelligence
Apple introduced a new product called Apple Intelligence in late 2024, which provides new OS-integrated AI capabilities across the Apple ecosystem. [^14] Apple Intelligence’s initial release was in October 2024, the ChatGPT integration was launched on December 11th, and further features (including Siri integration across third party apps) are coming soon [applepressrelease] [applepressrelease2-chatgpt]. Apple Intelligence will be able to process and perform AI tasks on E2EE content [applepressrelease]. According to its online documentation, Apple’s plan does not involve training any models on any “users’ private personal data or user interactions,” which presumably excludes E2EE content from use for training [appledeviceandservermodels].
In this section, we overview Apple’s plan for AI processing of E2EE content. We cover Apple’s plan not in order to spotlight Apple’s particular practices as opposed to those of other companies, but rather, simply because Apple’s plan is the best documented real-world example of AI processing of E2EE content on a major platform at the time of our writing. We use their plan as a convenient illustrative example for our exposition and analysis. Additional reasons, not focused primarily on Apple, that we find it compelling to study the example of Apple include: (1) in light of Apple’s proposal and the competitive AI landscape, it is likely that other major platforms are seriously considering similar programs; (2) in light of the large-scale weakening of encryption that could occur if AI assistant features in E2EE applications were deployed overly hastily, the kind of analysis in this paper is essential to conduct early; and (3) as other platforms consider similar programs, they may refer to Apple’s plan—conceivably both for technical reference and for normative and competitive justification for their own plans—as the earliest detailed documented proposal of the kind. We encourage readers to view the Apple examples throughout the paper as illustrative more than spotlighting, and to keep the broader landscape of other platforms in mind.
We commend Apple for their comparative transparency, which has facilitated our writing with reference to their concrete real-world example, while also acknowledging that despite that, their documentation leaves many critical questions unanswered (as discussed in Section 2.5.2). As we elaborate in later sections, we believe that stronger transparency and disclosure are essential when using AI to process E2EE content.
@@
Applications across the iOS ecosystem differ on the confidentiality guarantees they offer. Our work focuses only on data that is protected with E2EE, so in this section we explain what data in Apple devices is E2EE. There are two broad classes of applications for Apple devices: native applications, which are managed by Apple and come built-in with the device; and third-party applications, which are downloaded from the App Store. A few native applications are E2EE by default, such as iMessage, Facetime, and Keychain. However, storage on iCloud is not E2EE by default, so most native applications that rely on cloud storage are not E2EE. For additional security, users have the option of enabling Advanced Data Protection (ADP) [advanceddataprotection], a feature which enables E2EE for the majority of iCloud data, such as photos, notes, voice memos, etc. Some iCloud data is not E2EE even when ADP is activated, such as mail, contacts, and calendars. [^15] We refer readers to [advanceddataprotection] for a complete description of the confidentiality guarantees for various data types and native applications in iOS, both in the default case and when ADP is activated. In this section, we refer to native applications for which ADP enables E2EE as ADP-E2EE applications. In addition, ADP enables E2EE for certain Siri features, such as Siri shortcuts and other information (e.g., settings and personalization). It is not clear to us whether Siri transcripts are E2EE or not, even with ADP enabled.
The confidentiality guarantees of third-party applications vary by application. Some third-party applications are E2EE, such as Signal [signal], and WhatsApp [whatsapp].
@@
Siri and Apple Intelligence are able to process users’ content drawn from any application where the application developer has “integrated” Siri and Apple Intelligence functionalities [appleappintegration], and the user has not deactivated this integration in their settings [applesiriprivacy]. That is, if an application supports integration of Siri and Apple Intelligence, then integration/access is on by default unless a user turns it off in their settings.
@@
The second way in which Apple Intelligence may process E2EE data is through its content generation features, which may be accessible from within E2EE applications—this includes third-party E2EE and ADP-E2EE applications if ADP is enabled. For instance, image and emoji generation can be based on user-uploaded images and accessed through the iMessage app. If ADP is enabled, the generated images and emojis may be based on E2EE content.
2.5.1 Three-Tiered Approach
As users interact with Apple applications on their iPhones, iPads, or Macs, these applications may submit queries to the Apple Intelligence system. These queries, often a function of user data—and sometimes a function of user E2EE data—are then processed and responses are sent back to the user (e.g., producing short summaries of long text, or updating a contact’s information).
@@ 2.5.1 Three-Tiered Approach
Figure 4:
@@
Finally, Apple Intelligence sends certain user queries to ChatGPT. This feature is turned off by default, but can be manually enabled in settings [applesupport-chatgptext]. When the feature enabled, Siri examines each request and determines if ChatGPT “might be helpful to get the information” to respond to it [applesupport-siri]. The types of requests for which Siri may invoke ChatGPT may include: direct prompts to ChatGPT, text generation, and summarization [applesupport-siri]. If either ChatGPT is directly invoked or Siri deems ChatGPT could be “helpful” [applepressrelease], the user is asked in a pop-up prompt whether they would like to send the request to OpenAI for processing. [^20] For an example of what this looks like, see this example from the Apple Intelligence press release [applepressrelease]. By default, this prompt occurs every time a request could be sent to ChatGPT, but can be disabled in settings [^21] See footnote 16. [applesupport-siri]. Apple collects limited metadata about the queries sent to ChatGPT, including the number and size of requests [applesiriprivacy].
@@
2.5.2 What We Still Don’t Know
Based on available documentation at the time of writing, significant details remain unclear about Apple’s system. Open questions, to which the answers could be essential to independently evaluate the system’s security risks and compatibility with E2EE, include the following.
@@ 2.5.2 What We Still Don’t Know
    Will users be aware of each time Apple Intelligence is invoked?
@@
    Can users freely choose which Cases to allow? Could a user permit Case I (local processing)?
  •   $\bullet$
  •   $\bullet$
@@ 3 Our Definitions and Taxonomy
Built upon the technical background of the previous section, this section presents the definitions and taxonomy that underpin our findings.
@@ 3.1 Terminology
We write E2EE content or E2EE data to refer to any data which a service provider explicitly or implicitly indicates is protected by E2EE in an application, and any derivatives of such data. [^28] Derivatives include any mathematically correlated function of such data (excluding bit-length, as noted in footnote 7). This could be like the examples from Section 2.1.1, whether messages are happy/sad, or aggregate statistics about message content. Many claims of E2EE by service providers are not very specific regarding which data are E2EE—in such cases, we treat a generalized claim of E2EE in an application as an implicit claim that the primary application content is end-to-end encrypted. We do not treat non-primary application content as E2EE content unless a provider has explicitly claimed that that type of content is E2EE.
@@
3.2 A Taxonomy From “Strict E2EE” to “Not E2EE”
In this section, we present a taxonomy of applications ranging from “strict E2EE” to “not E2EE,” consisting of five categories in decreasing order of the degree to which they maintain the integrity of E2EE. As discussed in Section 2.1, E2EE applications today typically involve additional features and complexity beyond those explicitly modeled by formal definitions — and as such, they cover the full range of our first four categories, as further discussed below.
@@ 3.2 A Taxonomy From “Strict E2EE” to “Not E2EE”
No major E2EE application meets Category 1’s most stringent level of confidentiality. Certain commonplace features that enhance user experience involve the disclosure of (often small amounts of) E2EE content to third parties. [^29] One example, already discussed in Section LABEL:sec:background:e2ee-properties, is link previews. As such, even with these features, it has become common to refer to and advertise the application as “end-to-end encrypted,” both in industry and in academia, including among encryption experts. We refer to such features, which involve third-party processing of strictly limited types or quantities of plaintext-dependent content to provide a user-experience-enhancing feature, and which are commonplace in applications that provide the strongest E2EE guarantees in current practice, as caveat features. [^30] We recognize that term “strictly limited” is ambiguous here; we do not think it is realistic or helpful to articulate a precise threshold amount of processing data that would be acceptable. That said, we would not consider a feature involving third-party processing of a majority of E2EE content to be “strictly limited” processing. We would consider the caveat features that are commonplace in the most prominent E2EE applications in current practice to involve third-party processing that is “strictly limited” in the sense we mean here.
Category 2 (E2EE in current practice) encompasses a range of applications that do not qualify as “Strict E2EE” due to caveat features, but otherwise provide strong E2EE guarantees. To be in Category 2, no E2EE content may be seen or used by any third party at any point in the functioning of the application, except for caveat features, and applications must offer each user the option of disabling all caveat features in their own use of the application. (In other words, there must exist a configuration of application settings and features in which the application operates like a Category 1 application.) We place Signal, [^31] Showing the preview image of a link shared in Signal violates strict E2EE, for example.[signal2019link] and WhatsApp [^32] WhatsApp backup feature violates strict E2EE, for example.[whatsappbackup] in Category 2.
@@
Category 3 (E2EE with opt out) refers to applications which offer strong E2EE features as an option, but either these features are off by default, or the application allows users to opt out of these features. Note, in contrast, that Categories 1 and 2 definitionally preclude E2EE being off by default and/or users turning E2EE off. A Category 3 application’s E2EE features, when enabled, must be strong enough that no E2EE content may be seen or used by any third party at any point in the functioning of the application, except for caveat features. (That is, again, there must exist a configuration of settings and features in which the application operates like a Category 1 application.) We place Zoom [zoom] in Category 3, as it offers E2EE videoconferencing only when turned on in user settings. We consider Zoom to be in Category 3 even though a user can set E2EE as their personal default in their account settings (which means calls that that user sets up will be E2EE unless the user explicitly indicates otherwise); what we focus on is whether E2EE is off by default for any given user account. Notably, Category 3 applications lack the important systemic protections discussed in Section 2.1.2.
Category 4 (E2EE options available) encompasses applications that offer E2EE as an option for limited features, but does not offer any configuration where all primary application content is protected by E2EE. That is, Category 4 applications are not E2EE because they always involve processing of primary application content by third parties in the functioning of the application unless data is protected by E2EE as a narrow exception. Telegram allows users to turn on E2EE on a per-chat basis and it is always off by default[telegram]. At the time of writing, Meta’s Messenger is in Category 4, as it allows users to enable E2EE only in a special type of chat called “private chats”[messenger].
Finally, Category 5 (Not E2EE) describes applications that fall in none of the above categories. Category 5 does not preclude the use of encryption: for example, applications that use encryption but not end-to-end encryption fall in Category 5. As later sections elaborate, Category 5 applications are outside the scope of this paper—unless the service providers make claims that the application does offer E2EE.
4 Technical Implications of Integrating AI with E2EE
@@ 4.1 Security Considerations for AI assistants
As explained in Section 2, AI assistants interact with data during both training and inference. For an implementation of an AI assistant to be compatible with E2EE, it must uphold the guarantees of E2EE during both.

### v2

@@ 2.5.1 Apple Intelligence
Apple introduced a new product called Apple Intelligence in late 2024, which provides new OS-integrated AI capabilities across the Apple ecosystem. [^13] Apple Intelligence’s initial release was in October 2024; the ChatGPT integration was launched in December 2024 [applepressrelease, applepressrelease2-chatgpt]. Apple Intelligence is able to process and perform AI tasks on E2EE content [applepressrelease]. According to its online documentation, Apple’s plan does not involve training any models on any “users’ private personal data or user interactions,” which presumably excludes E2EE content from use for training [appledeviceandservermodels].
@@
Applications across the iOS ecosystem differ on the confidentiality guarantees they offer. Our work focuses only on data that is protected with E2EE, so in this section, we explain what data in Apple devices is E2EE. There are two broad classes of applications for Apple devices: native applications, which are managed by Apple and come built-in with the device; and third-party applications, which are downloaded from the App Store. A few native applications are E2EE by default, such as iMessage, Facetime, and Keychain. However, storage on iCloud is not E2EE by default, so most native applications that rely on cloud storage are not E2EE. For additional security, users have the option of enabling Advanced Data Protection (ADP) [advanceddataprotection], a feature that enables E2EE for the majority of iCloud data, such as photos, notes, voice memos, etc. Some iCloud data is not E2EE even when ADP is activated, such as mail, contacts, and calendars. [^14] We refer readers to [advanceddataprotection] for a complete description of the confidentiality guarantees for various data types and native applications in iOS, both in the default case and when ADP is activated. In this section, we refer to native applications for which ADP enables E2EE as ADP-E2EE applications. In addition, ADP enables E2EE for certain Siri features, such as Siri shortcuts and other information (e.g., settings and personalization). It is not clear to us whether Siri transcripts are E2EE or not, even with ADP enabled.
The confidentiality guarantees of third-party applications [^15] I.e., applications that run on Apple devices but are not developed by Apple. on Apple devices vary by application. Some third-party applications are E2EE, such as Signal [signal], and WhatsApp [whatsapp].
@@
Siri and Apple Intelligence will be able to process users’ content drawn from any application where the application developer has “integrated” Siri and Apple Intelligence functionalities [appleappintegration], and the user has not deactivated this integration in their settings [applesiriprivacy]. That is, if an application supports the integration of Siri and Apple Intelligence, then integration/access is on by default unless a user turns it off in their settings.
@@
The second way in which Apple Intelligence may process E2EE data is through its content generation features, which may be accessible from within E2EE applications—this includes third-party E2EE and ADP-E2EE applications if ADP is enabled. For instance, image and emoji generation can be based on user-uploaded images and accessed through the iMessage app. The generated images and emojis may be based on E2EE content if ADP is enabled.
Three-Tiered Approach
As users interact with Apple applications on their iPhones, iPads, or Macs, these applications may submit queries to the Apple Intelligence system. These queries, often a function of user data—and sometimes a function of user E2EE data—are then processed, and responses are sent back to the user (e.g., producing short summaries of long text or updating a contact’s information).
@@
Figure 4: Apple’s three-tiered approach
@@
Finally, Apple Intelligence sends certain user queries to ChatGPT. This feature is turned off by default, but can be manually enabled in settings [applesupport-chatgptext]. When the feature is enabled, Siri examines each request and determines if ChatGPT “might be helpful to get the information” to respond to it [applesupport-siri]. The types of requests for which Siri may invoke ChatGPT may include direct prompts to ChatGPT, text generation, and summarization [applesupport-siri]. If either ChatGPT is directly invoked or Siri deems ChatGPT could be “helpful” [applepressrelease], the user is asked in a pop-up prompt whether they would like to send the request to OpenAI for processing. [^20] For an example of what this looks like, see this example from the Apple Intelligence press release [applepressrelease]. By default, this prompt occurs every time a request could be sent to ChatGPT, but can be disabled in settings [^21] See footnote 16. [applesupport-siri]. Apple collects limited metadata about the queries sent to ChatGPT, including the number and size of requests [applesiriprivacy].
@@
What We Still Don’t Know
Based on the available documentation at the time of writing, significant details remain unclear about Apple’s system. Open questions, to which the answers could be essential to independently evaluate the system’s security risks and compatibility with E2EE, include the following.
@@
    Will users be aware each time Apple Intelligence is invoked?
  •   $\bullet$
@@
2.5.2 Samsung’s Galaxy AI
Samsung introduced a new series of AI-integrated phones in January 2025 [samsung-s25pressrelease]. Several new phones, such as the Galaxy S25 Ultra, Galaxy S25+, and Galaxy S25, will come with integrated AI assistants [samsung-s25pressrelease]: Samsung’s own Galaxy AI, some of whose features leverage Google’s Gemini [samsung-s25pressrelease]. While some previous Samsung phones were able to tap into certain Galaxy AI features, these newer models provide significant additional use cases.
What data on Samsung devices is E2EE?
E2EE applications on Samsung phones are either preinstalled on the device or can be downloaded from the Galaxy Store. Google Messages, for instance, comes with the phone and is E2EE by default between Google Messages users [samsung-defaultmessaging] [googlemessages-e2ee]. Cloud storage on Samsung devices is not E2EE by default. However, users have the option of enabling Enhanced Data Protection (EDP), a feature which enables E2EE for call log, message, clock, settings, and app backups [samsung-enhanceddataprotection] [^26] Large amounts of data such as message attachments may not be backed up under E2EE, even with Enhanced Data Protection [samsung-enhanceddataprotection].
Possible use cases involving E2EE data
Galaxy AI can process content in E2EE applications, including via the Gemini integration which can perform tasks across Google, Samsung, and third-party apps [samsung-s25pressrelease, samsung-galaxyaifaqs]. If users have Enhanced Data Protection enabled, Galaxy AI may also process additional content that is E2EE under EDP for purposes such as generating call transcripts and others [samsung-s25pressrelease].
Two-Tiered Approach
Galaxy AI queries are either processed on-device or outsourced to a cloud model. Samsung’s documentation notes that Google Messages features, “most translations”, and ambient wallpapers are processed on-device, whereas most other features, including auto-summarizing and image editing, are cloud-based [samsung-galaxyaifaqs].
On-device AI processes data locally with the Personal Data Engine, which is secured by hardware measures such as Knox Vault [samsung-knoxvault]. When processing on-device, queries and data are not used for training [samsung-dataprocessing-galaxyai].
According to Galaxy AI documentation as of January 2025, cloud models can use user query data for training [samsung-dataprocessing-galaxyai]. However, as explained at Samsung’s annual Galaxy Unpacked event in the same month, when Galaxy S25 series phones process queries in the cloud, user data is only used to respond to the query, is deleted as soon as the query has been responded to, and will not be used for training [samsung-galaxyunpacked]. Samsung partnered with Google Cloud in 2024 to provide Gemini-powered AI tools [samsung-googlecloud-partnership]. Several aspects of Galaxy AI’s cloud processing arrangement and security measures remain unclear, as further discussed below.
Galaxy AI is on by default on these newer phones, but some features require users to log in to their Samsung or Google accounts [samsung-galaxyaifaqs]. For example, Cross App Action (a feature that can perform tasks across different apps, including Google, Galaxy native, and third-party apps such as WhatsApp and Spotify) will require users to log into their Google account [samsung-aifeatures]. [^27] The Now Brief and Now Bar features, which provide users with relevant updates, will require users to log into their Samsung accounts [samsung-aifeatures]. And other features such as the Audio Eraser and AI Select also require a Samsung account [samsung-aifeatures]. Users have the options to disable cloud-based features while maintaining on-device processing capabilities, and to disable all AI features, in their settings [samsung-galaxyaifaqs].
What We Still Don’t Know
  • Does all Galaxy AI cloud processing take place on Google Cloud or just Gemini-based features?
  • What security measures are used for Galaxy AI’s cloud processing? How are encryption and trusted hardware employed, if at all?
  • What notices and disclosures are given to the user?
  • Are users aware of when AI is invoked?
  • Are users aware of how AI queries are being processed—locally or on the cloud? [^28] Users are able to select if they want to turn off AI features or turn off cloud-based AI features in their settings [samsung-galaxyaifaqs], but it remains unclear if users are told upfront how their queries are being processed.
2.5.3 Meta AI in WhatsApp
Meta features an integrated AI assistant in the E2EE messaging platform WhatsApp [whatsapp-faqmetaai]. This case, software integration instead of device integration, is slightly different than the previous two.
Meta AI can either be accessed in a standalone chat (between a user and the AI assistant) or it can join other conversations when it is summoned explicitly by a user sending a message containing the term “@MetaAI” [whatsapp-faqmetaai]. In the latter case, Meta AI is only able to read messages that mention @MetaAI or explicit replies to these messages [whatsapp-faqmetaai]. These boundaries are noted in the initial disclaimer [^29] See screenshots in Appendix B. the first time Meta AI is invoked, but do not appear afterward. Users are also given the option to delete AI messages by typing “/reset-all-ais” [whatsapp-faqusingmetaAI], though as noted below, it is not clear what exactly this does.
Meta AI always involves cloud processing; it does not run locally on user devices. Meta AI’s cloud processing infrastructure is not clear to us, although Meta has noted that the operation of Llama—on which Meta AI is based—is supported by “roster of partners … across the hardware and software ecosystem” including a number of third-party cloud providers [metaai-cloudmodels]. It is trained on a combination of publicly available online information, licensed information, and information shared on Meta products and services [metaai-faqs]. Queries sent from WhatsApp to Meta AI may be used for training, and this training data is kept by Meta “as long as [they] need it on a case-by-case basis” [metaai-faqs]; there is no option for users to opt out of this usage.
What We Still Don’t Know
  • Where does Meta AI’s cloud processing take place? How are Meta’s “partners … across the hardware and software ecosystem” involved, if at all?
  • What security measures, such as encryption and trusted hardware, are used for cloud processing?
  • What exactly happens when users “reset” and delete chats with AIs?
@@ 3.1 Terminology
We write E2EE content or E2EE data to refer to any data which a service provider explicitly or implicitly indicates is protected by E2EE in an application, and any derivatives of such data. [^32] Derivatives include any mathematically correlated function of such data (excluding bit-length). Many claims of E2EE by service providers are not very specific regarding which data are E2EE—in such cases, we treat a generalized claim of E2EE in an application as an implicit claim that the primary application content is end-to-end encrypted. We do not treat non-primary application content as E2EE content unless a provider has explicitly claimed that that type of content is E2EE.
@@
3.2 A Taxonomy From “Strict E2EE” to “No E2EE”
In this section, we present a taxonomy of applications ranging from “Strict E2EE” to “No E2EE,” consisting of five categories in decreasing order of the degree to which they maintain the integrity of E2EE. As discussed in Section 2.1, modern E2EE applications typically involve additional features and complexity beyond those explicitly modeled by formal definitions. As such, they cover the full range of our first four categories.
@@ 3.2 A Taxonomy From “Strict E2EE” to “No E2EE”
No major E2EE application meets Category 1’s most stringent level of confidentiality, at the time of writing. Certain commonplace features that enhance user experience involve the disclosure of (often small amounts of) E2EE content to third parties. [^33] One example, already discussed in Section 2.1.5, is link previews. As such, even with these features, it has become common to refer to and advertise the application as “end-to-end encrypted,” both in industry and in academia, including among encryption experts. We refer to such features, which involve third-party processing of strictly limited types or quantities of plaintext-dependent content to provide a user-experience-enhancing feature, and which are commonplace in applications that provide the strongest E2EE guarantees in current practice, as caveat features. [^34] We recognize that term “strictly limited” is ambiguous here; we do not think it is realistic or helpful to articulate a precise threshold amount of processing data that would be acceptable. That said, we would not consider a feature involving third-party processing of a majority of E2EE content to be “strictly limited” processing. We would consider the caveat features that are commonplace in the most prominent E2EE applications in current practice to involve third-party processing that is “strictly limited” in the sense we mean here.
Category 2 (E2EE in current practice) encompasses a range of applications that do not qualify as “Strict E2EE” due to caveat features, but otherwise provide strong E2EE guarantees. To be in Category 2, no E2EE content may be seen or used by any third party at any point in the functioning of the application, except for caveat features, and applications must offer each user the option of disabling all caveat features in their own use of the application. (In other words, there must exist a configuration of application settings and features in which the application operates like a Category 1 application.) We place Signal [^35] Showing the preview image of a link shared in Signal violates strict E2EE, for example[signal2019link]. and WhatsApp [^36] WhatsApp’s (non-E2EE) backup feature violates strict E2EE, for example[whatsappbackup]. in Category 2.
@@
Category 3 (E2EE with opt out) refers to applications which offer strong E2EE features as an option, but either these features are off by default, or the application allows users to opt out of these features. Note, in contrast, that Categories 1 and 2 definitionally preclude E2EE being off by default and/or users turning E2EE off. A Category 3 application’s E2EE features, when enabled, must be strong enough that no E2EE content may be seen or used by any third party at any point in the functioning of the application, except for caveat features. (That is, again, there must exist a configuration of settings and features in which the application operates like a Category 1 application.) We place Zoom [zoom] in Category 3, as it offers E2EE videoconferencing only when turned on in user settings. We consider Zoom to be in Category 3 since a user can set E2EE as their personal default in their account settings (which means calls that that user sets up will be E2EE unless the user explicitly indicates otherwise); what we focus on is whether E2EE is off by default for any given user account. Notably, Category 3 applications lack the important systemic protections discussed in Section 2.1.3.
Category 4 (E2EE options available) encompasses applications that offer E2EE as an option for limited features, but does not offer any configuration where all primary application content is protected by E2EE. That is, Category 4 applications are not E2EE because they always involve processing of primary application content by third parties in the functioning of the application unless data is protected by E2EE as a narrow exception. Telegram allows users to turn on E2EE on a per-chat basis and it is always off by default [telegram].
Finally, Category 5 (No E2EE) describes applications that fall in none of the above categories. Category 5 does not preclude the use of encryption: for example, applications that use encryption but not end-to-end encryption fall in Category 5. As later sections elaborate, Category 5 applications are outside the scope of this paper—unless the service providers make claims that the application does offer E2EE.
4 Security Implications of Integrating AI with E2EE
@@ 4.1 Security Considerations for AI assistants
As explained in Section 2, AI assistants interact with data during both training and inference. For an implementation of an AI assistant to be compatible with E2EE, it must uphold the guarantees of E2EE during both training and inference.